### v1

# Dynamical breaking of inversion symmetry and strong second harmonic generation with nonlinear phonons

Abstract
We show how crystalline inversion symmetry can be dynamically broken by optical phonons with generic, hardening Kerr-like non-linearities. The symmetry-broken state is reached through a parametric instability that can be accessed by driving close to half the phonon resonance. After the onset of the instability, the system settles to a steady state with inversion-symmetry breaking phonon trajectories and strong second harmonic generation. The time averaged positions of the atoms are displaced relative to equilibrium in the steady state, resulting in a rectification of the driving signal.
Introduction
The non-equilibrium behavior of condensed matter systems currently attracts considerable attention due to its potential for on-demand control over materials (basov2017_on_demand, ; bloch2022strongly_corr_el_photon, ; rudner2020band, ). Out-of-equilibrium phonons are of particular interest, as lattice distortions have an immediate impact on the electronic properties of solids (forst2011nonlinear_phononics, ; mankowsky2016nonlinear_phononics_2, ). Nonlinear phonon resonances can be expoited to create, enhace and manipulate superconductivity (mankowsky2014_enhanced_superconductivity, ; knap2016phonon_superconduct_theory, ; babadi2017phonon_superconduct_theory, ; cavalleri2018photo_induced_supercond, ; liu2020_res_phonon_light_induced_superc, ), magnetism (fechner2018magnetophononics_review, ; afanasiev2021ultrafast_control_magnetic_phonons, ; disa2023_phonon_ferromag_induce, ; luo2025terahertz_control_magno_phononics, ) and other states of matter (nova2019phonon_ferroelectricity, ; ning2023_phonon_induced_hidden_quadrupolar, ; kaplan2025_spatiotemporal_phonons, ; kaplan2025spatiotemporal_phonons_first_principles, ).
Here, we show how nonlinear phonons can be driven in unconventional ways to create symmetry-forbidden electromagnetic responses and lattice deformations that dynamically break the underlying crystal symmetries. We focus on inversion symmetry, which is known to prohibit second harmonic generation (SHG) and rectification (Boyd_nonlinear_optics, ). We demonstrate however that, even for a very generic third order nonlinearity, this rule can be circumvented by a taylored driving protocol. For chiral phonons carrying finite angular momentum (juraschek2025chiral_phonons_rev, )(Note1, ), we demonstrate how the strong second harmonic generation leads to diverse Lissajous like non-inversion-symmetric phonon trajectories. Such trajectories create structured magnetic fields that can influence the dynamics of electrons and spins on a microscopic level (yaniv2025multicolor, ; luo2023large_magnetic_chiral_phonons, ; juraschek2022giant_magn_field_chiral_phonons, ; xiong2022effective_magn_field_chiral, ).
A model for nonlinear chiral phonons
We will discuss both, degenerate chiral optical phonons, for which the resonance frequency does not depend on the sense of motion (kahana2024_chiral_rectification, ; cheng2020_dirac_semi_phonon_zeeman, ; mustafa2025origin_phonon_zeeman_mos2, ), as well as phonons with split frequencies for right- and left-handed motion (zhang2015chiral_phonons_hexagonal, ; zhu2018observation_chiral_phonons, ; ishito2023truly_chiral_phonons, ; ueda2023chiral_phonons_xrays, ). We begin with the former case. A simple model for nonlinear, degenerate chiral phonons is given by the Hamiltonian (kahana2024_chiral_rectification, )
$H=P_{x}^{2}+P_{y}^{2}+\frac{\Omega_{0}^{2}}{2}\left(Q_{x}^{2}+Q_{y}^{2}\right)+\frac{\beta}{4}\left(Q_{x}^{2}+Q_{y}^{2}\right)^{2}+V_{\mathrm{l-m}}.$ (1)
[…]
$V_{\mathrm{l-m}}=-\mathbf{E}\cdot\left(\mathbf{p}_{x}+\mathbf{p}_{y}\right),$ (2)
where the electric dipole moments of the phonon components are given by $\mathbf{p}_{n}=\mathbf{Z}_{n}Q_{n}$, with effective electric charges $\mathbf{Z}_{n}$. For simplicity, we assume $\mathbf{Z}_{n}\propto\hat{\mathbf{e}}_{n}$. Then, for circularly polarized light, the phonon equations of motion read
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}+\beta Q_{x}\left(Q_{x}^{2}+Q_{y}^{2}\right)$ $=Z_{x}E_{x}\cos\omega t$ (3)
$\ddot{Q}_{y}+2\gamma\dot{Q}_{y}+\Omega_{0}^{2}Q_{y}+\beta Q_{y}\left(Q_{x}^{2}+Q_{y}^{2}\right)$ $=Z_{y}E_{y}\sin\omega t,$ (4)
where we included a damping term with damping rate $\gamma$.
Instability for a single phonon component
To show how the symmetry breaking instability emerges, we first focus on a single phonon component $Q_{x}\left(t\right)$ and set $E_{y}=0$. It is useful to divide $Q_{x}\left(t\right)$ into parts composed of odd and even harmonics:
$Q_{i}\left(t\right)=Q_{i,\mathrm{odd}}\left(t\right)+Q_{i,\mathrm{even}}\left(t\right),$ (5)
which are, respectively, antisymmetric and symmetric under a time translation by half the oscillation period of the electromagnetic field:
$Q_{i,\mathrm{odd}}\left(t+\frac{\pi}{\omega}\right)$ $=-Q_{i,\mathrm{odd}}\left(t\right),$
$Q_{i,\mathrm{even}}\left(t+\frac{\pi}{\omega}\right)$ $=Q_{i,\mathrm{even}}\left(t\right).$ (6)
Naively, one expects $Q_{i,\mathrm{even}}$ to vanish, because the odd-order nonlinearity in the equations of motion (4) does not couple even harmonics to the driving. However, in the following, we describe a route create even harmonics via a parametric instability.
Let us first study the onset of this instability. Since we assume, for the moment, that $E_{y}=0$, we can set $Q_{y}=0$ and consider the dynamics for the $Q_{x}$ mode alone. Using the decomposition of Eq. (5), we can separate the equation (3) into equations for $Q_{x,\mathrm{odd}}$ and $Q_{x,\mathrm{even}}$. We use that, at the onset of the instability, $Q_{i,\mathrm{even}}$ will be very small, such that $\left|Q_{i,\mathrm{even}}\right|\ll\left|Q_{i,\mathrm{odd}}\right|$. Then, the equation for the odd part, neglecting contributions stemming from $Q_{i,\mathrm{even}}$, reads
$\ddot{Q}_{x,\mathrm{odd}}+2\gamma\dot{Q}_{x,\mathrm{odd}}+\Omega_{0}^{2}Q_{x,\mathrm{odd}}+\alpha Q_{x,\mathrm{odd}}^{3}=Z_{x}E_{x}\cos\left(\omega t\right).$ (7)
This is the equation of a simple driven Duffing oscillator. For our purposes it is sufficient to approximate the response $Q_{x,\mathrm{odd}}$ with the fundamental harmonic and write
$Q_{x,\mathrm{odd}}\approx F_{x}\left(E_{x}\right)\cos\left(\omega t+\varphi_{x}\right).$ (8)
$F_{x}\left(E_{x}\right)$ is then found by inverting the amplitude equation
$F_{x}^{2}\left[4\gamma^{2}\omega^{2}+\left(\left(\omega^{2}-\Omega_{0}^{2}\right)-\frac{3}{4}\beta F_{x}^{2}\right)^{2}\right]=Z_{x}^{2}E_{x}^{2}.$ (9)
For the even component $Q_{x,\mathrm{even}}$, we find the Mathieu equation
$\ddot{Q}_{x,\mathrm{even}}+2\gamma\dot{Q}_{x,\mathrm{even}}$
$+\tilde{\Omega}_{0}^{2}\left(E_{x}\right)\left[1+h\left(E_{x}\right)\cos\left(2\omega t+2\phi_{x}\right)\right]Q_{x,\mathrm{even}}=0$ (10)
where $\tilde{\Omega}_{0}\left(E_{x}\right)$ is an effective, amplitude dependent resonance frequency [see Fig. 1a)] given by
$\tilde{\Omega}_{0}\left(F_{x}\right)=\Omega_{0}\sqrt{1+\frac{3\alpha}{2\Omega_{0}^{2}}F_{x}^{2}\left(E_{x}\right)},$ (11)
and $h\left(E_{x}\right)=3\alpha F_{x}^{2}\left(E_{x}\right)/\left[2\tilde{\Omega}_{0}^{2}\left(E_{x}\right)\right]$. We used Eq. (8) to approximate $Q_{x,\mathrm{odd}}^{2}$. It is then the constant-in-time part of $Q_{x,\mathrm{odd}}^{2}$ that modifies the resonance frequency of the mode and leads to a blue shift, while the oscillating part of $Q_{x,\mathrm{odd}}^{2}$ acts as a parametric driving for $Q_{x,\mathrm{even}}$. The Mathieu equation (14) is known to exhibit parametric instabilities for $\tilde{\Omega}_{0}\left(E_{x}\right)=n\omega$, with $n$ a positive integer (Landau_Lifshitz_Mechanics, ). However, $Q_{x,\mathrm{even}}$has to obey Eq. (6), which excludes the $n=1$ resonance. The $n=2$ resonance, however, is allowed, and leads to the symmetry-breaking instability we want to study. Here $\tilde{\Omega}_{0}\left(E_{x}\right)=2\omega$, such that for driving slightly above half the original resonance frequency of $\Omega_{0}$, we expect a response at $\tilde{\Omega}_{0}\left(E_{x}\right)$ – i.e., we expect strong SHG.
As is typical for parametric resonances, the instability occurs in a small frequency window where for $\Delta=2\omega-\tilde{\Omega}_{0}$ holds (see e.g. (turyn1993damped_Mathieu, ), p. 394 (Note2, ))
$\frac{\tilde{\Omega}_{0}}{24}\left(3\left(\frac{4\gamma^{2}}{\tilde{\Omega}_{0}^{2}}-\sqrt{h^{4}-\frac{64\gamma^{2}}{\tilde{\Omega}_{0}^{2}}}\right)+2h^{2}\right)<\Delta$
$<\frac{1}{24}\tilde{\Omega}_{0}\left(3\left(\frac{4\gamma^{2}}{\tilde{\Omega}_{0}^{2}}+\sqrt{h^{4}-\frac{64\gamma^{2}}{\tilde{\Omega}_{0}^{2}}}\right)+2h^{2}\right).$ (12)
Here, and in what follows, we omit writing out the $E_{x}$-dependence of $F_{x}$, $\tilde{\Omega}_{0}$, $h$ and $\Delta$ explicitly, except when it is needed for clarity. The frequency blue shift and the instability window are illustrated with the results of a numerical simulation in Fig. 1 a).
To overcome damping effects, a minimal driving amplitude is required. This threshold amplitude $E_{x,*}$ can be calculated by setting $\Delta=0$. To leading order in $\gamma/\Omega$, we find $h\left(E_{x,*}\right)=\sqrt{8\gamma/\Omega_{0}}$ where $\Omega_{0}$ is the $E_{x}$-independent resonance frequency of Eq. (1). This expression can be inverted for $E_{x}$ using Eq. (9). For small damping, the threshold electric field amplitude is then given by
$E_{x,*}\approx\frac{\sqrt{3}\Omega^{3}}{2^{3/4}\sqrt{\beta}Z_{x}}\left(\frac{\gamma}{\Omega}\right)^{1/4}$ (13)
As can be expected, $E_{x,*}$ is lowered by a strong nonlinearity $\beta$ and increased by a larger $\gamma$. The result of Eq. (13) is confirmed by numerical simulations as shown in Fig. 1 c).
Upon going through the parametric instability at $2\omega\approx\tilde{\Omega}_{0}\left(E_{x}\right)$, the phonons reach a stable trajectory. This behavior is not uncommon in nonlinear systems (dykman1998fluctuational_transitions_parametric_osc, ; marthaler2006parametric_osc_quantum_switching, ), however, in our case the steady state is characterized by strong fundamental and second harmonic response, as well as a considerable DC offset. The spectrum of $Q_{x}$ in this steady state, obtained by solving Eq. (3) numerically with $E_{y}=0$, is shown in Fig. 1 b). We note in passing, that the instability and steady state studied here are known in nonlinear systems literature, although the only extensive study to our knowledge is presented in Ref. (ys1991_duffing_symm_breaking_SHG, ). Below we extend our results to the chiral system of Eqs. (3) and (4).
[…]
DC offset
We now show that the instability and steady state outlined above, necessarily imply the presence of a static displacement of the atoms from their equilibrium positions. To see this, we averagy Eq. (3) over one period of the drive. As above, we assume $E_{y}=0$ and therefore $Q_{y}=0$. Writing $Q_{x}\left(t\right)=\sum_{a}Q_{x,a}\cos\left(n\omega t+\varphi_{x,a}\right)$, and truncating the series at $n=2$, find
$\omega_{0}^{2}Q_{x,0}+\beta\left[\frac{3}{2}\left(Q_{x,1}^{2}+Q_{x,2}^{2}\right)Q_{x,0}\right$.
$\quad+\left.\frac{3}{4}Q_{x,1}^{2}Q_{x,2}\cos\left(2\varphi_{x,1}-\varphi_{x,2}\right)+Q_{x,0}^{3}\right].$
This equation has one non-trivial, real solution for $Q_{x,0}$. To leading order in $Q_{x,1}$ and $Q_{x,2}$, it reads
$Q_{x,0}=-\frac{3\beta}{4\omega_{0}^{2}}Q_{x,1}^{2}Q_{x,2},$
showing that any response at the second harmonic is accompanied by a DC offset. Being third order in the first and second harmonic amplitudes, we expect the DC offset to be smaller in magnitude, it can however, still be sizable (see Fig. (1) b). We conclude that although the symmetry breaking instability is triggered by an oscillating driving field, inversion symmetry is still broken on average. This will result in constant in time electric field produced by the dipoles $\mathbf{p}=Z_{x}Q_{x,0}$, where $Z_{x}$ is the effective electric charge of the phonon mode in question, i.e. the driving signal is rectified.
Exploiting resonant modes
We note that phonons with frequencies close to $\Omega_{0}/2$ can be exploited to resonantly enhance the otherwise off-resonant driving. Consider an auxiliary, IR active phonon mode $P_{A}$, such that it couples to $Q_{x}$ via a term
$H_{PQ}=\lambda Q_{x}P_{A}.$
This coupling preserves the original inversion invariance of the system and leads to $\lambda P_{A}$ taking over the role of the electric field in Eq. (3). $P_{A}$ can then be a regular IR active mode following (to zeroth order in $\lambda$) the equation of motion $\ddot{P}_{A}+2\gamma_{A}\dot{P}_{A}+\omega_{A}P_{A}=Z_{A}E_{x}\cos\left(\omega_{A}t\right)$, such that it accumulated the energy of the electric field over a number of $\sim\omega_{A}/\gamma_{A}$ cycles. Due to the inherent nonlinear blue-shift of the effective resonance frequency $\tilde{\Omega}_{0}$ [Eq. (11)], the driving power can be adjusted such that the resonance frequency $\omega_{A}$ of the auxiliary phonon mode $P_{A}$ exactly hits $\omega_{R}=\tilde{\Omega}_{0}/2$, very similar to the situation depicted in Fig. 1 a), where, the driving frequency is fixed to $0.6\Omega_{0}$, while the driving amplitude is increased. Around $E_{x}=0.4\,\text{\AA }\sqrt{u}\Omega_{0}^{2}/Z_{x}$, the resonance condition is fulfilled and the systems enters the symmetry breaking regime. Thus the auxiliary mode does not have to be located at exactly half the resonant frequency of the $Q_{x}$ mode, rather, the blue shift can be exploited to access the instability at the twice the frequency of the auxiliary mode by adjusting the driving strength.
Collective instability of the x and y modes
Having studied the symmetry breaking instability for a single phonon component driven by linearly polarized light, we now turn to the full chiral system consisting of modes $Q_{x}$ and $Q_{y}$ described by Eqs. (3), (4). Numerically, we observe that the instability intervals are larger if $E_{x}\neq E_{y}$, i.e. the driving electromagnetic field is eliptically polarized. To rationalize this observation, we perform a stability analysis for the two-component equations (3), (4) following Ref. (Landau_Lifshitz_Mechanics, ).
We first derive the two-component analogue of Eq. (10), which is given by
$\ddot{Q}_{x,\mathrm{even}}+2\gamma\dot{Q}_{x,\mathrm{even}}+\left[\tilde{\Omega}_{x}^{2}+\frac{\alpha}{2}\left(3F_{x}^{2}-F_{y}^{2}\right)\cos\left(2\omega t\right)\right]Q_{x,\mathrm{even}}+\alpha F_{x}F_{y}\sin\left(2\omega t\right)Q_{y,\mathrm{even}}$ $=0$
$\ddot{Q}_{y,\mathrm{even}}+2\gamma\dot{Q}_{y,\mathrm{even}}+\left[\tilde{\Omega}_{y}^{2}-\frac{\alpha}{2}\left(3F_{y}^{2}-F_{x}^{2}\right)\cos\left(2\omega t\right)\right]Q_{y,\mathrm{even}}+\alpha F_{x}F_{y}\sin\left(2\omega t\right)Q_{x,\mathrm{even}}$ $=0,$ (14)
with $\tilde{\Omega}_{x/y}^{2}=\Omega_{x/y}^{2}+\alpha\left(3F_{x/y}^{2}+F_{y/x}^{2}\right)/2$, where $F_{y}$ is defined analogously to $F_{x}$ in Eq. (8). As above, we expect parametric resonances near $2\omega=\tilde{\Omega}_{x/y}$, where $\tilde{\Omega}_{x}\neq\tilde{\Omega}_{y}$ for $F_{x}\neq F_{y}$. For now, let us choose the case $2\omega=\tilde{\Omega}_{x}$, such that the instability occurs for the $Q_{x}\left(t\right)$ component.
The oscillating terms in Eqs. (14) couple harmonics with frequencies $2\omega$, $4\omega$, … and DC terms. For the stability analysis, we therefore choose the ansatz
$Q_{i}$ $=a_{i,1}\sin\left(2\omega t\right)+a_{i,2}\sin\left(4\omega t\right)$
$\quad+b_{i,0}+b_{i,1}\cos\left(2\omega t\right)+b_{i,2}\cos\left(4\omega t\right).$ (15)
Furthermore, we neglect $\gamma$ for the duration of this analysis. While $\gamma$ determines the instability threshold amplitudes of the electromagnetic fields [see Eq. (13)], its effects become less important for driving amplitudes above the threshold, i.e., for any driving amplitude, $\gamma$ can be always chosen small enough that our analysis is accurate. We again search for the instability window for the detuning $\Delta=2\omega-\tilde{\Omega}_{x}$, such that the mode amplitudes $a_{i,n}$ and $b_{i,n}$ grow exponentially for $\Delta_{\mathrm{min}}<\Delta<\Delta_{\mathrm{max}}$. At $\Delta=\Delta_{\mathrm{max}/\mathrm{min}}$, the amplitudes will be constant. The boundaries of the instability interval $\Delta_{\mathrm{max}/\mathrm{min}}$ are then found by inserting the ansatz (15) into Eqs. (14) and assuming that $a_{i,n}$ and $b_{i,n}$ are indeed constant (Landau_Lifshitz_Mechanics, ). After a lengthy calculation, in which we compare the coefficients of different harmonics after inserting the ansatz (15) into Eqs. (14), we find that, to fourth order in $F_{x}$ and $F_{y}$
$\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ $=F_{y}^{4}\left(\frac{\alpha^{2}}{16\Omega_{0}^{3}}+\frac{287\alpha^{4}F_{x}^{4}}{576\Omega_{0}^{7}}+\frac{47\alpha^{3}F_{x}^{2}}{192\Omega_{0}^{5}}\right)$
$\quad+F_{y}^{2}\left(\frac{77\alpha^{3}F_{x}^{4}}{192\Omega_{0}^{5}}-\frac{5\alpha^{2}F_{x}^{2}}{8\Omega_{0}^{3}}\right)+\frac{9\alpha^{2}F_{x}^{4}}{16\Omega_{0}^{3}}.$ (16)
The full result is too long to be quoted here but is easily found using computer algebra. Eq. (16) is valid for small $F_{x}$ and $F_{y}$, i.e. for small driving amplitudes. It is interesting to study the behavior of $\Delta$ close to $F_{y}=F_{x}$, i.e. for nearly perfect circular polarization. Writing $F_{y}=F_{x}+F_{\epsilon}$, we find
$\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ $=-F_{\epsilon}^{3}\left(\frac{9\Omega_{0}}{8F_{x}^{3}}+\frac{\alpha^{2}F_{x}}{\Omega_{0}^{3}}+\frac{51\alpha}{16F_{x}\Omega_{0}}\right).$ (17)
Notice that for $F_{y}\geq F_{x}$ (we choose both amplitudes positive w.l.o.g.), we have $\Delta_{\mathrm{max}}\leq\Delta_{\mathrm{min}}$, which indicates that the system is stable. For $F_{y}>F_{x}$, the $Q_{x}$ and $Q_{y}$ components switch places, and the instability occurs for $2\omega=\tilde{\Omega}_{y}$. The analysis for this case is completely analogous with $F_{x}$ and $F_{y}$ , as well as $\Omega_{x}$ and $\Omega_{y}$ interchanged. We therefore conclude that, in general, the instability occurs either for the $Q_{x}$ or the $Q_{y}$ component, depending on whether $F_{x}$ or $F_{y}$ is larger.
Figure 2: The instability window $\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ [see the discussion above Eq. (16)] as a function of $F_{y}/F_{x}$ is plotted. The full solution for the ansatz of Eq. (15) is plotted as a blue solid line in the main figure and in the inset. The dashed red line indicates the approximation of Eq. (16). The dashed green line in the inset shows the approximation of Eq. (17), which is valid close to $F_{y}=F_{x}$. Stable regions, where $\Delta_{\mathrm{max}}\leq\Delta_{\mathrm{min}}$ are marked yellow [see discussion below Eq. (17)]. We conclude that the system is stable at $F_{x}=F_{y}$ which holds for driving with perfectly circularly polarized light, i.e. $E_{x}=E_{y}$. Some amount of ellipticity of the driving electromagnetic field is necessary to access the instability.
The expansions of Eqs. (16) and (17) in comparison to the full result are shown in Fig. 2 for small amplitudes $F_{x}$, $F_{y}$. If driven with perfectly circularly polarized light with $E_{x}=E_{y}$, resulting in $F_{x}=F_{y}$, the time-dependent terms in Eqs. (14) cancel each other, and we find $\Delta_{\mathrm{min}}=\Delta_{\mathrm{max}}$; the symmetry breaking state cannot be reached. Thus, to induce the symmetry breaking transition, elliptically polarized light must be used (Note3, ).
In Fig. 1 d), we plot the numerical solution for $Q_{x}\left(t\right)$, $Q_{y}\left(t\right)$ for elliptically polarized light with $E_{y}=0.25E_{x}$. The resulting Lissajous trajectory breaks the inversion symmetry of the Hamiltonian (1), due to the large second harmonic component of $Q_{x}\left(t\right)$.
Beyond the instability at $\omega\approx\Omega_{0}/2$ that we have studies so far, higher order, inversion symmetry breaking instabilities at frequencies $\omega\approx\Omega_{0}/2n$ can be induced. These instabilities generate higher order even harmonics. In general, the required threshold driving powers are larger for higher order parametric instabilities, and grow according to $E_{x,*}^{2}\sim\gamma^{1/\left(2n\right)}$ (Landau_Lifshitz_Mechanics, ). Fig. 1 e) shows the phonon trajectories for the $n=2$ instability, where $Q_{x}\left(t\right)$ exhibits a strong fourth harmonic component.
Non-degenerate chiral modes
Finally, we investigate dynamical symmetry breaking for non-degenerate chiral Phonons. A toy-model with split frequencies for phonons of opposite chiralities is obtained by substituting $P_{i}\rightarrow P_{i}-\kappa A_{i}$ in the Hamiltonian of Eq. (1). Here $\mathbf{A}=B_{\mathrm{eff}}\left[-Q_{y},Q_{x},0\right]$ takes the role of an effective magnetic vector potential acting on the motion of the phonon components. To linear order in $\kappa$, the above substitution is equivalent to adding the term $\kappa\mathbf{B}_{\mathrm{eff}}\cdot\mathbf{L}$ to the Hamiltonian (1), i.e.
$H\rightarrow H+\kappa\mathbf{B}_{\mathrm{eff}}\cdot\mathbf{L},$ (18)
where $\mathbf{L}=\left(Q_{x}P_{y}-Q_{y}P_{x}\right)\hat{\mathbf{e}}_{z}$ is the phonon angular momentum and $\mathbf{B}_{\mathrm{eff}}=\left[0,0,B_{\mathrm{eff}}\right]$. Solving the linearized equations of motion, we find the phonon eigenfrequencies
$\Omega_{0,\pm}=\Omega_{0}\pm\kappa B_{\mathrm{eff}},$ (19)
where the $\pm$-signs corresponds to right- and left-handed motion, respectively.
We find that the symmetry-breaking instability described above can also be achieved with non-degenerate chiral phonons. The right- and left-handed modes can be accessed separatly, depending on the polarization of the driving electromagnetic field. Because of the driving-induced blue-shift, the instabilty can be accessed for the two non-degenerate modes at the same frequency, but at different driving powers. In agreement with the results for degenerate chiral phonons discussed above [see Eq. (17)], we find that perfectly circular polarized light is ineffective in inducing the symmetry breaking. A certain amount of ellipticity is necessary, to trigger the instability at half the resonance frequency. We present these results in Fig. 3.
Figure 3: Symmetry breaking with non-degenerate chiral phonons [see Eq. (18)]. The phonon frequencies are split according to Eq. (19): $\Omega_{+}$ corresponds to right-handed motion, while $\Omega_{-}$ corresponds to a left-handed rotation. The two modes are accessed with light of opposite polarizations, fitting their respective sense of motion. To excite the $\Omega_{+}$ mode, we use $\mathbf{E}=E_{0}\left[-\left(1-\delta\right)\cos\omega t,\sin\omega t,0\right]$, and for the $\Omega_{-}$ mode, $\mathbf{E}=E_{0}\left[\cos\omega t,\left(1-\delta\right)\sin\omega t,0\right]$, with $\delta=0.25$. As for degenerate chiral phonons, a slight detuning from circularity $\delta$ is necessary, in order to trigger the instability at half the resonance frequency [see Eq. (17)]. The figure combines the results of two runs, in which the two chiralities were simulated separately. The resonance frequencies´ exhibit a driving amplitude dependent blue-shift, such that the instability occurs at different powers, for the two chiralities.
Conclusion
In conclusion, we have described a new, symmetry breaking steady state for driven chiral, nonlinear phonons. This state is characterized by strong second harmonic generation and by the emergence of non-oscillating electric fields. These effects, being forbidden by the inversion symmetry of the underlying lattice, can serve as sharp experimental signatures of inversion symmetry breaking in the steady state and the effects presented in this manuscript. Beyond possible applications for second harmonic generation and rectification, the study of interactions of chiral phonons in the newly described state with electrons and other collective modes (e.g. magnons (kahana2024_chiral_rectification, )) offers an intriguing avenue for uncovering novel out-of-equilibrium correlated states and dynamical phase transitions.
Acknowledgements.
We acknowledge comments by the anonymous referee of one of our previous publications (kiselev2019squid_spectroscopy, ), which, eventually, led to the research reported here, as well as useful discussions with Mark Rudner. We also thank Jonas F. Karcher who motivated us to work on this manuscript, by pointing out that “a Nature article about chiral phonons was on his newsfeed, and chiral phonons seem to be a hot topic” (juraschek2025chiral_phonons_rev, ). This project received funding from the Horizon Europe Marie Skłodowska-Curie Action program under Grant Agreement 101155351.