### v3

# Dynamical breaking of inversion symmetry, strong second harmonic generation, and ferroelectricity with nonlinear phonons

Abstract
We show how crystalline inversion symmetry can be dynamically broken by optical phonons with generic, hardening Kerr-like non-linearities. The symmetry-broken state is reached through a parametric instability that can be accessed by driving close to half the phonon frequency. The system then settles to a steady state with inversion-symmetry breaking phonon trajectories and strong second harmonic generation. The time averaged positions of the atoms are displaced relative to equilibrium, resulting in a ferroelectric rectification of the driving signal. For circularly polarized phonons, complex Lissajous-like trajectories can be achieved.
Introduction
The non-equilibrium behavior of condensed matter systems currently attracts considerable attention due to its potential for on-demand control over materials (basov2017_on_demand, ; bloch2022strongly_corr_el_photon, ; rudner2020band, ). Out-of-equilibrium phonons are of particular interest, as lattice distortions have an immediate impact on the electronic properties of solids (forst2011nonlinear_phononics, ; mankowsky2016nonlinear_phononics_2, ). Nonlinear lattice dynamics can be expoited to enhace and manipulate superconductivity (mankowsky2014_enhanced_superconductivity, ; knap2016phonon_superconduct_theory, ; babadi2017phonon_superconduct_theory, ; cavalleri2018photo_induced_supercond, ; liu2020_res_phonon_light_induced_superc, ), magnetism (fechner2018magnetophononics_review, ; afanasiev2021ultrafast_control_magnetic_phonons, ; disa2023_phonon_ferromag_induce, ; luo2025terahertz_control_magno_phononics, ) and other states of matter (nova2019phonon_ferroelectricity, ; ning2023_phonon_induced_hidden_quadrupolar, ; kaplan2025_spatiotemporal_phonons, ; kaplan2025spatiotemporal_phonons_first_principles, ). Other effects predicted for driven, nonlinear collective modes include non-trivial steady states (kiselev2024MFPD_plasmons_nat, ; kiselev2024_PRB_MFPD, ; kiselev2025exciting, ; wanic2025entanglement, ; kaplan2025_spatiotemporal_phonons, ; kaplan2025spatiotemporal_phonons_first_principles, ; kaplan2025optically_induced_goldstone_faraday, ), chaos (wanic2024magnetoelectric, ), and parametric amplification (cartella2018parametric, ; buzzi2021higgs_lasing, ; michael2024photonic, ).
Here, we show how IR-active phonons oscillating in anharmonic interatomic potentials can be driven in unconventional ways to break underlying crystal symmetries and create symmetry-forbidden electromagnetic responses. We focus on inversion symmetry, which is known to prohibit second harmonic generation (SHG) and rectification (Boyd_nonlinear_optics, ). We demonstrate that, for a very generic hardening Kerr-like nonlinearity, this rule can be circumvented by a taylored driving protocol, resulting in a symmetry-breaking steady state with strong phononic SHG and DC responses. We stress that no double well-structure with metastable ferroelectric states, as e.g. exploited in Ref. (li2019_driven_metastable_ferroelectricity, ), is required.
The DC response leads to a static displacement of atoms from their equilibrium positions, generating a constant-in-time dipolar electric field. This behavior is analogous to that of an equilibrium ferroelectric, where – below a critical temperature – dipolar fields form due to spontaneous, symmetry-breaking lattice deformations. In our case, the static dipolar response is triggered by a dynamical breaking of inversion symmetry, and we argue that the predicted effect can be used to create an on-demand ferroelectric. In particular, we show that the non-equilibrium ferroelectric response exhibits hysteretic behavior, which underlines that the effect is not merely a nonlinear piezoelectric response (disa2020polarizing_symmetry_breaking_light_nonlinearity, ), but instead a symmetry-breaking out-of-equilibrium phase. We show that this phase is robust to noise, and is accessible at reasonable field strenghts.
We then explore the above effects in a system of circularly polarized phonons (juraschek2025chiral_phonons_rev, ), which are sometimes referred to as chiral phonons (kahana2024_chiral_rectification, ), and have been used to manipulate magnetic states. Here, the strong SHG leads to Lissajous-like, non-inversion-symmetric phonon trajectories. Such trajectories create structured magnetic fields influencing the dynamics of electrons and spins on a microscopic level (luo2023large_magnetic_chiral_phonons, ; juraschek2022giant_magn_field_chiral_phonons, ; xiong2022effective_magn_field_chiral, ), and provide an interesting way of engineering electronic behavior (yaniv2025multicolor, ).
Symmetry-breaking instability and second harmonic generation
Figure 1: a) Phonons [Eq. (18)] driven with a linearly polarized electric field $E_{x}\cos\left(\omega t\right)$ oscillating at frequency $\omega=0.6\Omega_{0}$. The effective phonon frequency $\tilde{\Omega}_{0}\left(E_{x}\right)$, given in Eq. (10), exhibits a blue shift as $E_{x}$ is increased. Around $\tilde{\Omega}_{0}\left(E_{x}\right)\approx 2\omega$, the system enters the symmetry breaking state with strong second harmonic generation. In the symmetry-breaking regime, the resonance curve $\tilde{\Omega}_{0}\left(E_{x}\right)$ is interrupted. A small damping of $\gamma=0.001$ was chosen to have a large instability interval for clarity. b) Spectrum of $Q_{x}\left(t\right)$ in the symmetry breaking state. c) Amplitude ratio of second and first harmonics across the symmetry breaking transition for different dampings $\gamma$. d) Hysteresis of the mean polarization $p_{x}=Z_{x}F_{x,\text{0}}=\left(2\pi/\omega\right)\int_{0}^{2\pi/\omega}dt\,Q_{x}\left(t\right)$ when a static bias field $E_{\mathrm{\mathrm{DC}}}$ is swept through positive and negative values in a triangle-wave-like manner.
To show how the symmetry breaking instability emerges, we first focus on a single IR-active phonon mode $Q_{x}\left(t\right)$. For a generic, inversion symmetric system, the dynamics of $Q_{x}\left(t\right)$ is governed by the Hamiltonian
$H=\frac{1}{2}P_{x}^{2}+\frac{\Omega_{0}^{2}}{2}Q_{x}^{2}+\frac{\beta}{4}Q_{x}^{4}+V_{\mathrm{l-m}}.$ (1)
[…]
$V_{\mathrm{l-m}}=-\mathbf{E}\cdot\mathbf{p}_{x},$ (2)
where the electric dipole moment of the phonon is given by $\mathbf{p}_{x}=\mathbf{Z}_{x}Q_{x}$, with effective charge $\mathbf{Z}_{x}$. We assume $\mathbf{Z}_{x}\propto\hat{\mathbf{e}}_{x}$. Then, for circularly polarized light, the phonon equation of motion reads
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}+\beta Q_{x}^{3}$ $=Z_{x}E_{x}\cos\omega t$ (3)
where we included a damping term with damping rate $\gamma$.
It is often assumed that the third order nonlinearity of Eq. (3) does not generate even harmonics of the driving frequency $\omega$. Indeed, shifting the time coordinate according to $t\rightarrow t+\frac{\pi}{\omega}$ changes the sign of the right hand side, which seems to enforce the symmetry $Q_{x}\left(t+\frac{\pi}{\omega}\right)=-Q_{x}\left(t\right)$, excluding any response at frequencies $2n\omega$, where $n\in\mathbb{N}$, including zero. We now show, that this symmetry, stemming from the inversion symmetry of the Hamiltonian (1), can be dynamically broken.
It is useful to divide $Q_{x}\left(t\right)$ into parts composed of odd and even harmonics:
$Q_{x}\left(t\right)=Q_{x,\mathrm{odd}}\left(t\right)+Q_{x,\mathrm{even}}\left(t\right),$ (4)
which are, respectively, antisymmetric and symmetric under a time translation by half the oscillation period of the electromagnetic field:
$Q_{x,\mathrm{odd}}\left(t+\frac{\pi}{\omega}\right)$ $=-Q_{x,\mathrm{odd}}\left(t\right),$
$Q_{x,\mathrm{even}}\left(t+\frac{\pi}{\omega}\right)$ $=Q_{x,\mathrm{even}}\left(t\right).$ (5)
We now show how even harmonics ($Q_{i,\mathrm{even}}$) can be created via a parametric instability. Let us first study the onset of this instability. Using the decomposition of Eq. (4), we can separate Eq. (A 7) into cupled equations for $Q_{x,\mathrm{odd}}$ and $Q_{x,\mathrm{even}}$. We use that, at the onset of the instability, $Q_{i,\mathrm{even}}$ will be very small, such that $\left|Q_{i,\mathrm{even}}\right|\ll\left|Q_{i,\mathrm{odd}}\right|$. Then, the equation for the odd part, neglecting contributions stemming from $Q_{i,\mathrm{even}}$, reads
$\ddot{Q}_{x,\mathrm{odd}}+2\gamma\dot{Q}_{x,\mathrm{odd}}+\Omega_{0}^{2}Q_{x,\mathrm{odd}}+\alpha Q_{x,\mathrm{odd}}^{3}=Z_{x}E_{x}\cos\left(\omega t\right).$ (6)
For our purposes it is sufficient to approximate the response $Q_{x,\mathrm{odd}}$ with the fundamental harmonic and write
$Q_{x,\mathrm{odd}}\approx F_{x,1}\left(E_{x}\right)\cos\left(\omega t+\varphi_{x}\right).$ (7)
$F_{x,1}\left(E_{x}\right)$ is then found by inverting the amplitude equation
$F_{x,1}^{2}\left[4\gamma^{2}\omega^{2}+\left(\left(\omega^{2}-\Omega_{0}^{2}\right)-\frac{3}{4}\beta F_{x,1}^{2}\right)^{2}\right]=Z_{x}^{2}E_{x}^{2}.$ (8)
For the even component $Q_{x,\mathrm{even}}$, we find the Mathieu equation
$\ddot{Q}_{x,\mathrm{even}}+2\gamma\dot{Q}_{x,\mathrm{even}}$
$+\tilde{\Omega}_{0}^{2}\left(E_{x}\right)\left[1+h\left(E_{x}\right)\cos\left(2\omega t+2\varphi_{x}\right)\right]Q_{x,\mathrm{even}}=0$ (9)
where $\tilde{\Omega}_{0}\left(E_{x}\right)$ is an effective, amplitude dependent resonance frequency [see Fig. 1a)] given by
$\tilde{\Omega}_{0}\left(E_{x}\right)=\Omega_{0}\sqrt{1+\frac{3\beta}{2\Omega_{0}^{2}}F_{x,1}^{2}\left(E_{x}\right)},$ (10)
and $h\left(E_{x}\right)=3\alpha F_{x,1}^{2}\left(E_{x}\right)/\left[2\tilde{\Omega}_{0}^{2}\left(E_{x}\right)\right]$. We used Eq. (7) to approximate $Q_{x,\mathrm{odd}}^{2}$. It is then the constant-in-time part of $Q_{x,\mathrm{odd}}^{2}$ that modifies the resonance frequency of the mode and leads to a blue shift, while the oscillating part of $Q_{x,\mathrm{odd}}^{2}$ acts as a parametric driving for $Q_{x,\mathrm{even}}$. The Mathieu equation (A 8) is known to exhibit parametric instabilities for $\tilde{\Omega}_{0}\left(E_{x}\right)=n\omega$, with $n$ a positive integer (Landau_Lifshitz_Mechanics, ). However, $Q_{x,\mathrm{even}}$has to obey Eq. (5), which excludes the $n=1$ resonance. The $n=2$ resonance, however, is allowed, and leads to the symmetry-breaking instability we want to study. Here $\tilde{\Omega}_{0}\left(E_{x}\right)=2\omega$, such that for driving slightly above half the original resonance frequency of $\Omega_{0}$, we expect a response at $\tilde{\Omega}_{0}\left(E_{x}\right)$ – i.e., we expect strong SHG.
As is typical for parametric resonances, the instability occurs in a small frequency window where for $\Delta=2\omega-\tilde{\Omega}_{0}$ holds (see e.g. (turyn1993damped_Mathieu, ), p. 394 (Note2, ))
$\frac{\tilde{\Omega}_{0}}{24}\left(3\left(\frac{4\gamma^{2}}{\tilde{\Omega}_{0}^{2}}-\sqrt{h^{4}-\frac{64\gamma^{2}}{\tilde{\Omega}_{0}^{2}}}\right)+2h^{2}\right)<\Delta$
$<\frac{1}{24}\tilde{\Omega}_{0}\left(3\left(\frac{4\gamma^{2}}{\tilde{\Omega}_{0}^{2}}+\sqrt{h^{4}-\frac{64\gamma^{2}}{\tilde{\Omega}_{0}^{2}}}\right)+2h^{2}\right).$ (11)
Here, and in what follows, we omit writing out the $E_{x}$-dependence of $F_{x,1}$, $\tilde{\Omega}_{0}$, $h$ and $\Delta$ explicitly, except when it is needed for clarity. The blue shift and the instability window are shown with the results of a numerical simulation in Fig. 1 a) (see AppendixAppendix.1.1 for a note on the initial conditions used). We stress that, while in our analytical calculation we approximated the odd response by a single harmonic and linearized in $Q_{x,\mathrm{even}}$, we used the full Eq. (3) for the simulation. As can be expected there are discrepancies due to the approximations made, however, the instability window is clearly identifiable.
To overcome damping effects, a minimal driving amplitude is required. This threshold amplitude $E_{x,*}$ can be calculated by setting $\Delta=0$. To leading order in $\gamma/\Omega$, we find $h\left(E_{x,*}\right)=\sqrt{8\gamma/\Omega_{0}}$. This expression can be inverted for $E_{x}$ using Eq. (8). For small damping, the threshold electric field amplitude is then given by
$E_{x,*}\approx\frac{\sqrt{3}\Omega^{3}}{2^{3/4}\sqrt{\beta}Z_{x}}\left(\frac{\gamma}{\Omega_{0}}\right)^{1/4}.$ (12)
Notice that $\Omega_{0}$ is the bare resonance frequency, but the blue-shift has been taken into account in the derivation of Eq. (12). As can be expected, $E_{x,*}$ is lowered by a strong nonlinearity $\beta$ and increased by a larger $\gamma$. The result of Eq. (12) is confirmed by numerical simulations as shown in Fig. 1 c). The threshold fields for different (bare) phonon resonance frequencies $\Omega_{0}$ are given in Table 1. Since $E_{x,*}\sim\Omega_{0}^{3}$ holds, much stronger amplitudes are required at higher frequencies. The amplitudes are, however, in an experimentally accessible range.
$\Omega_{0}/2\pi$ $1\,\mathrm{THz}$ $3\,\mathrm{THz}$ $5\,\mathrm{THz}$
$E_{x,*}$ $1.5\cdot 10^{6}\,\mathrm{V/m}$ $4.1\cdot 10^{7}\,\mathrm{V/m}$ $1.9\cdot 10^{8}\,\mathrm{V/m}$
[…]
Symmetry breaking steady-state and pulsed excitation
The above analysis demonstrates that the usual solution, consisting of a strong response at the fundamental harmonic and perturbative corrections at odd harmonics, is unstable if driving frequency and strength are tuned such that the condition of Eq. (11) at $2\omega\approx\tilde{\Omega}_{0}\left(E_{x}\right)$ is fulfilled. This instability is associated with strong second harmonic generation due to the parametric resonance in Eq. (A 8). Investigating the dynamics of the nonlinear phonon mode in this regime, we find that the system converges to a stable steady-state which breaks the inversion symmetry of the system. Besides a strong second harmonic, this symmetry-breaking is characterized by a relatively large DC offset. The spectrum of $Q_{x}$ in the steady state, obtained by solving Eq. (3) numerically, is shown in Fig. 1 b). We note in passing, that the instability and steady state studied here have been described in nonlinear systems literature, although the only extensive study, to our knowledge, is presented in Ref. (ys1991_duffing_symm_breaking_SHG, ).
For low frequency phonons the instability is accessible already at relatively small driving fields (see Table1), which might be produced by continuous wave lasers. However, reaching the symmetry breaking with pulsed signals is also possible. We performed a series of numerical experiments using the same parameters as in the continuous wave case, and find that short pulses with as few as five oscillation cycles (see Appendix Appendix.1) are sufficient to observe both second harmonic generation and a DC offset. However, flat-top pulses must be used, because, due to the nonlinear blue-shift, the frequency and amplitude of the drive must be tuned toghether. The dashed lines in Fig.1 b) show the response spectrum obtained for a flat top pulse of five optical cycles. The finite duration of the pulse introduces spectral broadening, however, the SHG and the DC offset are clearly visible.
Before extending our results to a system of chiral phonons, we investigate how the DC component of the steady-state at $2\omega\approx\tilde{\Omega}_{0}\left(E_{x}\right)$ leads to a ferroelectric response, and show how auxiliary phonon modes can be exploited to trigger the symmetry breaking instability resonantly.
Ferroelectricity, hysteresis, stability to noise
We now show that the symmetry-breaking steady-state outlined above, necessarily implies the presence of a static displacement of the atoms from their equilibrium positions. To see this, we average Eq. (3) over one period of the drive. Writing $Q_{x}\left(t\right)=\sum_{a}F_{x,a}\cos\left(a\omega t+\varphi_{x,a}\right)$, and truncating the series at $a=2$, we find
$\omega_{0}^{2}F_{x,0}+\beta\left[\frac{3}{2}\left(F_{x,1}^{2}+F_{x,2}^{2}\right)F_{x,0}\right$.
$\quad+\left.\frac{3}{4}F_{x,1}^{2}F_{x,2}\cos\left(2\varphi_{x,1}-\varphi_{x,2}\right)+F_{x,0}^{3}\right].$ (13)
This equation has one non-trivial, real solution for $F_{x,0}$. To leading order in $F_{x,1}$ and $F_{x,2}$, it reads
$F_{x,0}=-\frac{3\beta}{4\omega_{0}^{2}}F_{x,1}^{2}F_{x,2},$ (14)
showing that any response at the second harmonic is accompanied by a DC offset. Being third order in the first and second harmonic amplitudes, we expect the DC offset to be smaller in magnitude, it can however, still be sizable [see Fig. (1) b)].
We conclude that although the symmetry breaking instability is triggered by an oscillating driving field, inversion symmetry is statically broken. This results in constant-in-time electric fields produced by the dipoles $\mathbf{p}=\mathbf{Z}_{x}F_{x,0}$, where $\mathbf{Z}_{x}$ is the effective electric charge of the phonon mode in question. The driving signal is thus rectified and triggers a ferroelectric response.
This dynamically triggered ferroelectric response shares significant similarities with equilibrium ferroelectrics, showing that it can be used as a road to on-demand light driven ferroelectricity. To substantiate this claim, we demonstrate that hysteresis can be observed when the phonon mode is in the symmetry-breaking state. To this end we add a static component $E_{\mathrm{DC}}$ to the driving field in Eq. (3) and sweep it through positive and negative values in a triangle-wave-like manner. We numerically integrate Eq. (3) while slowly changing $E_{\mathrm{DC}}$. Fig.1d) shows the mean polarization
$p_{x}=Z_{x}F_{x,\text{0}}=\left(2\pi/\omega\right)\int_{0}^{2\pi/\omega}dt\,Q_{x}\left(t\right),$ (15)
which exhibits a hysteresis. Interestingly, the hysteresis curve is pinched – also a feature known from equilibrium ferroelectrics (xu2016pinched_hysteresis_ferroelectrics, ).
[…]
Figure 2: Collapse of the symmetry-breaking steady state under the influence of noise. In the case of thermal noise and $\Omega_{0}/2\pi=3\,\mathrm{THz}$, the critical noise strength $A\approx 0.1Z_{x}E_{x}$ corresponds to a temperature of $\sim 700\,\mathrm{K}$. The same parameters as in Fig.1 were used.
Exploiting auxiliary IR- and Raman-active modes
Auxiliary modes can be used to further simplify reaching the symmetry-breaking steady state. Infrared-active phonons with frequencies $\omega_{A}$ close to $\Omega_{0}/2$ can be exploited to resonantly enhance the otherwise off-resonant driving: consider a phonon mode $Q_{A}$ with a coupling term
$H_{qQ}=\lambda Q_{x}Q_{A}.$ (16)
This coupling preserves the original inversion symmetry of the system and leads to $\lambda Q_{A}$ taking over the role of the electric field in Eq. (A 7). The mode $q_{A}$ can be pumped resonantly, such that the energy of the drive is stored over $\sim\omega_{A}/\gamma_{A}$ cycles, resulting in a large amplitude oscillation of $Q_{A}$. Due to the nonlinear blue-shift of $\tilde{\Omega}_{0}$ [Eq. (10)], the frequency of the auxiliary phonon $\omega_{A}$ can be adjusted by choosing a suitable driving amplitude, such that it exactly hits $\omega_{A}=\tilde{\Omega}_{0}/2$.
Raman-active modes can be used in a similar way. Here the lowest order, inversion symmetric coupling Hamiltonian is given by
$H=AQ_{R}Q_{x}^{2}.$ (17)
$Q_{R}$ is an auxiliary Raman mode. As shown in Appendix Appendix.3, including the coupling of Eq. (17) again leads to an equation of motion for $Q_{x}$ that can be mapped onto the original Eq. (A 7). Thus Raman active modes can, too, be used to drive the system into the symmetry-breaking state. This might me advantageous, since driving these modes allows for greater optical penetration depths compared to infrared-active modes.
Nonlinear, circularly polarized phonons
Figure 3: The Lissajous trajectory of phonon coordinates $Q_{x}\left(t\right)$ and $Q_{y}\left(t\right)$ when driven into the symmetry-breaking state using elliptically polarized light with $\mathbf{E}=E_{x}\left[\cos\left(\omega t\right),0.25\sin\left(\omega t\right)\right]$. The inversion symmetry of Eq. (18) is broken dynamically.
We now investigate the effects of dynamical symmetry breaking on circularly polarized phonons. While sometimes such phonons are referred to as chiral (kahana2024_chiral_rectification, ), this terminology has been questioned (juraschek2025chiral_phonons_rev, ). We will discuss both, degenerate circularly polarized optical phonons (kahana2024_chiral_rectification, ; cheng2020_dirac_semi_phonon_zeeman, ; mustafa2025origin_phonon_zeeman_mos2, ), as well as phonons with split frequencies for right- and left-handed motion (zhang2015chiral_phonons_hexagonal, ; zhu2018observation_chiral_phonons, ; ishito2023truly_chiral_phonons, ; ueda2023chiral_phonons_xrays, ).
A simple model for nonlinear, degenerate, circularly polarized phonons is given by the Hamiltonian (kahana2024_chiral_rectification, )
$H=\frac{P_{x}^{2}}{2}+\frac{P_{y}^{2}}{2}+\frac{\Omega_{0}^{2}}{2}\left(Q_{x}^{2}+Q_{y}^{2}\right)+\frac{\beta}{4}\left(Q_{x}^{2}+Q_{y}^{2}\right)^{2}+V_{\mathrm{l-m}}.$ (18)
Here, $Q_{x}$ and $Q_{y}$ are the coordinates of two orthogonal phonon modes. The light-matter coupling $V_{\mathrm{l-m}}$ is given by
$V_{\mathrm{l-m}}=-\mathbf{E}\cdot\left(\mathbf{p}_{x}+\mathbf{p}_{y}\right),$ (19)
where $\mathbf{p}_{n}=\mathbf{Z}_{n}Q_{n}$ are the electric dipole moments of the phonon components, with effective charges $\mathbf{Z}_{n}$. For simplicity, we assume $\mathbf{Z}_{n}\propto\hat{\mathbf{e}}_{n}$. The equations of motion for elliptically polarized light are given in AppendixAppendix.4, where we added damping at rate $\gamma$.
Solving the equations of motion (A 7) numerically, we observe that the instability intervals are larger for elliptical polarization with $E_{x}\neq E_{y}$. A stability analysis for the system following Ref. (Landau_Lifshitz_Mechanics, ), for which we refer to AppendixAppendix.4, rationalizes this observation.
The result of a numerical simulation of the phonon trajectory in the symmetry-breaking state is shown in in Fig. 3 for driving with elliptically polarized light where $E_{y}=0.25E_{x}$. The resulting Lissajous trajectory breaks the inversion symmetry of the Hamiltonian (18), due to the large second harmonic component of $Q_{x}\left(t\right)$.
We note that beyond the instability at $\omega\approx\Omega_{0}/2$, higher order, inversion symmetry breaking instabilities at frequencies $\omega\approx\Omega_{0}/n$, where $n$ is an even number can be induced. These generate more complex Lissajous-figures, however the required threshold driving amplitudes grow according to $E_{x,*}\sim\gamma^{1/2n}$ (Landau_Lifshitz_Mechanics, ).
Finally, we investigate dynamical symmetry breaking for non-degenerate circularly polarized phonons (zhang2015chiral_phonons_hexagonal, ) (also called “false chiral phonons” (juraschek2025chiral_phonons_rev, )). Such phonons can appear in systems with broken time-reversal symmetry (bonini2023CrI3_chiral_phonons, ) and have been observed in monolayer TMDs (zhu2018observation_chiral_phonons, ). A toy-model with split frequencies for phonons of opposite chiralities, is obtained by substituting $P_{i}\rightarrow P_{i}-\kappa A_{i}$ in the Hamiltonian of Eq. (18) (chen2019chiral_chen_review, ). Here $\mathbf{A}=B_{\mathrm{eff}}\left[-Q_{y},Q_{x},0\right]$ takes the role of an effective magnetic vector potential. To linear order in $\kappa$, the above substitution is equivalent to adding the term $\kappa\mathbf{B}_{\mathrm{eff}}\cdot\mathbf{L}$ to the Hamiltonian (18), where $\mathbf{L}=\left(Q_{x}P_{y}-Q_{y}P_{x}\right)\hat{\mathbf{e}}_{z}$ is the phonon angular momentum and $\mathbf{B}_{\mathrm{eff}}=\left[0,0,B_{\mathrm{eff}}\right]$. Solving the linearized equations of motion, we find the phonon eigenfrequencies
$\Omega_{0,\pm}=\Omega_{0}\pm\kappa B_{\mathrm{eff}},$ (20)
where the $\pm$-signs correspond to right- and left-handed motion, respectively.
We find that the symmetry-breaking instability described above can also be achieved with non-degenerate chiral phonons and present these results in Fig. 4.
Figure 4: Symmetry breaking with non-degenerate chiral phonons [see Eq. (20)]. The phonon frequencies are split according to Eq. (20): $\Omega_{\pm}$ corresponds to right/left-handed motion. The two modes are accessed with light of opposite polarizations. To excite the $\Omega_{+}$ mode, we use $\mathbf{E}=E_{0}\left[-\left(1-\delta\right)\cos\omega t,\sin\omega t,0\right]$, and for the $\Omega_{-}$ mode, $\mathbf{E}=E_{0}\left[\cos\omega t,\left(1-\delta\right)\sin\omega t,0\right]$, with $\delta=0.25$ and $\omega=0.62\Omega$. As for degenerate chiral phonons, a slight detuning from circularity $\delta$ is necessary, in order to trigger the instability at half the resonance frequency [see Eq. (A 11)]. The figure combines the results of two runs, in which the two chiralities were simulated separately. The resonance frequencies exhibit a driving amplitude dependent blue-shift, such that the instability occurs at different powers, for the two chiralities.
Conclusion
We have described a new, inversion-symmetry-breaking steady state for driven nonlinear phonons. This state is characterized by strong second harmonic generation and by the emergence of ferroelectric behavior. These responses – forbidden by the inversion symmetry of the underlying lattice – can serve as sharp experimental signatures of the physics presented in this manuscript.
Ferroelectrics above the transition (ginzburg1949theory_ferroelectrics, ; cochran1959_cochran_ferroelectrics1, ; cochran1960_cochran_ferroelectrics2, ) are natural platforms for realizing the effects. The softening phonon modes responsible for the transition will have low frequencies and strong anharmonicity (scott1974soft_modes_review, ; pal2021origin_thz_soft_modes, ; alyabyeva2021_barium_hexaferrite_soft_modes, ), resulting in low threshold powers $E_{x,*}$ as shown in Eq. (12).
The symmetry-breaking state can be induced with flat-top pulses as short as five optical cycles, and at reachable field strengths (Table1). Beyond possible applications for second harmonic generation, rectification, and driven, on-demand ferroelectricity, the study of interactions of phonons in the newly described state with electrons and collective modes offers a route to novel out-of-equilibrium states.
Acknowledgements.
We acknowledge comments by the anonymous referee of one of our previous publications (kiselev2019squid_spectroscopy, ), which, eventually, led to the research reported here, as well as useful discussions with Mark Rudner. We also thank Jonas F. Karcher who motivated us to work on this manuscript, by pointing out that “a Nature article about chiral phonons was on his newsfeed (juraschek2025chiral_phonons_rev, ), and chiral phonons seem to be a hot topic”. This project received funding from the Horizon Europe Marie Skłodowska-Curie Action program under Grant Agreement 101155351.
Appendix
Appendix.1 Achieving symmetry breaking with pulsed excitations
We perform simulations to show the symmetry-breaking steady state can be reached with short pulsed signals. We confine the oscillating electric field in Eq. (3) to a window of length $T_{p}=5\cdot 2\pi/\omega$. This mimics a flat-top pulse (see Fig.A 1) whose amplitude is constant throughout $T_{p}$. The constant amplitude is necessary, because due to the blue-shift of $\tilde{\Omega}_{0}\left(E_{x}\right)$ both amplitude and frequency of the drive have to be matched with each other (see Fig. 1). Fig. 1 shows the flat-top pulse and the response of the phonon mode $Q_{x}\left(t\right)$.
Figure A 1: Reaching the symmetry-breaking state is possible with a short flat-top pulse of only five optical cycles. The upper plot shows the driving electric field, while the lower plot shows the response with visible second harmonic and DC offset components. When the pulse ends, the system switches to unforced oscillations at frequency $\Omega_{0}$. The of $Q_{x}\left(t\right)$ is show in Fig.1 of the main text.
Appendix.1.1 A note on initial conditions and symmetry breaking
We started our simulation runs at $t=0$ and used the forcing term $E_{x}\left(t\right)=E_{0,x}\cos\left(\omega t\right)$, effectively this corresponds to $E_{x}\left(t\right)\sim\Theta\left(t\right)\cos\left(\omega t\right)$. The response of Eq. (3) to this forcing necessarily consists of particular solution oscillating at $\omega$ and persisting for $t\rightarrow\infty$, and a homogeneous solution oscillating at $\tilde{\Omega}_{0}$ that fixes the boundary conditions at $t=0$, and decays due to damping. The homogeneous solution has both finite $Q_{\mathrm{even}}$ and $Q_{\mathrm{odd}}$ components since $\tilde{\Omega}_{0}$ is not necessarily a multiple of $\omega$. It thus initializes the parametric instability in Eq. (9), which gives $Q_{\mathrm{even}}\left(t\right)=0$ for $Q_{\mathrm{even}}\left(0\right)=0$ and $\dot{Q}_{\mathrm{even}}\left(0\right)=0$, acting as an initial forcing for $Q_{\mathrm{even}}$, and determining the sign of the symmetry breaking.
Appendix.2 Noise robustness
In our noise simulations, $\eta\left(t\right)$ is a normally-distributed random function on a time grid with time points $t_{i}$ separated by intervals $\Delta t=5\cdot 10^{-3}/\Omega_{0}$. For simplicity, we use cubic splines to interpolate $\eta\left(t\right)$ between the $t_{i}$, and use a standard numerical integrator (SciPy’s odeint). This does not influence the physical results since $\Delta t\ll 2\pi/\Omega_{0}$. The symmetry-breaking steady state is stable up to a threshold noise amplitude $A_{*}$.
To relate the noise amplitude $A_{*}$ to a temperature, we calculate the internal energy of the oscillator at finite temperatures via $U=-\partial_{\tilde{\beta}}\log Z$:
$U$ $=-\frac{1}{2\pi\hbar}\frac{\partial}{\partial\left(1/k_{B}T\right)}\log\int dq\int dpe^{-\frac{H\left(p,q\right)}{k_{B}T}}$
$=\frac{1}{8}\left(\frac{\Omega^{4}K_{\frac{5}{4}}\left(\frac{\Omega^{4}}{8\beta k_{B}T}\right)}{\beta K_{\frac{1}{4}}\left(\frac{\Omega^{4}}{8\beta k_{B}T}\right)}-\frac{\Omega^{4}}{\beta}+2k_{B}T\right)$
$\approx k_{B}T-\frac{3\beta\left(k_{B}T\right)^{2}}{4\Omega^{4}}-\mathcal{O}\left(\beta^{2}\right).$
Assuming a phonon resonance frequency of 33\,THz, and extracting the internal energy $U$ corresponding to noise of amplitude $A_{*}$ (in absence of driving), we estimate that $A_{*}$ corresponds to a temperature of about 700700\,K.
Appendix.3 Pumping via Raman-active modes
In centrosymmetric materials, Raman- and IR-active modes are mutually exclusive. They can be, however, nonlinearly coupled to each other via the coupling Hamiltonian of Eq. (17), where $Q_{R}$ is the Raman active and $Q_{x}$ the IR active mode. The equations of motion for the coupled system can then be written down as
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}+\beta Q_{x}^{3}$ $=2AQ_{x}Q_{R}$
$+Z_{x}E_{0}\cos\left(\omega t\right)$ (A 1)
$\ddot{Q}_{R}+2\gamma\dot{Q}_{R}+\Omega_{R}^{2}Q_{R}+\beta_{R,2}Q_{R}^{2}+\beta_{R,3}Q_{R}^{3}.$
$\qquad=AQ_{x}^{2}+\frac{1}{2}\chi_{R}E_{0}^{2}\left(1+\cos\left(2\omega t\right)\right)$ (A 2)
We assume that the drive is in resonance with the Raman-active mode, such that $2\omega\approx\Omega_{R}$. Furthermore, we assume that $\Omega_{0}$ is distinct from $\Omega_{R}$, such that the driving is off-resonant with $\Omega_{0}$, and the direct, off-resonant driving of $Q_{x}$ via $Z_{x}$ is neglectable. We will furthermore be interested in instabilites of $Q_{x}$ induced by the Raman mode. At the instability onset, which we want to study here, $Q_{x}$ will be small, and we can neglect higher order terms in this variable. The simplified system of equations reads
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}-2AQ_{x}Q_{R}$ $=0$ (A 3)
$\ddot{Q}_{R}+2\gamma\dot{Q}_{R}+\Omega_{R}^{2}Q_{R}+\beta_{R,2}Q_{R}^{2}+\beta_{R,3}Q_{R}^{3}$
$\qquad=\frac{1}{2}\chi_{R}E_{0}^{2}\left(1+\cos\left(2\omega t\right)\right).$ (A 4)
Inverting the second equation, we find, neglecting higher harmonics
$Q_{R}\approx F_{R,0}\left(E_{0}\right)+F_{R,2}\left(E_{0}\right)\cos\left(2\omega t\right),$ (A 5)
where $F_{R,0}$ and $F_{R,2}$ are the amplitudes of DC and second harmonic (with respect to the driving frequency) components. Inserting this into the first equation gives
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\tilde{\Omega}_{0}\left(E_{0}\right)\left[1-h\left(E_{0}\right)\cos\left(2\omega t\right)\right]Q_{x}=0.$ (A 6)
Here $\tilde{\Omega}_{0}\left(E_{0}\right)=\Omega_{0}^{2}+F_{R,0}\left(E_{0}\right)$ is the effective, blue shifted resonance frequency of the IR active mode and $h\left(E_{0}\right)=2AF_{R,2}\left(E_{0}\right)$. This equation maps onto Eq. () of the main text. Showing, that a Raman-active mode can be used to trigger the symmetry-breaking instability of the IR active mode. The frequency of the resonant Raman-active mode $\Omega_{R}\approx 2\omega$ has to be reasonbly close to $\Omega_{0}$. However, some detuning is possible, because the effective resonance frequency $\tilde{\Omega}_{0}\left(E_{0}\right)$ experiences a blue shift. This blue shift can be exploted to adjust the effective resonance frequency by adjusting the driving amplitude $E_{0}$.
Similarly, an auxiliary IR-active mode with a frequency close to half the resonance $\Omega_{0}$ can be used, instead of the Raman-active mode. An advantage of using an auxiliary (Raman or IR-active mode) is that the mode will story the energy supplied by the drive over many cycles. Thus the threshold amplitude for the instability might be easier reach.
Figure A 2: Illustration of the instability condition of Eq. (11) for the same parameteres as in Fig. 1a).
Appendix.4 Collective instability of the x and y modes
The equations of motion (with the addition of a damping term) following form Eq. (18) read
$\ddot{Q}_{x}+2\gamma\dot{Q}_{x}+\Omega_{0}^{2}Q_{x}+\beta Q_{x}\left(Q_{x}^{2}+Q_{y}^{2}\right)$ $=Z_{x}E_{x}\cos\omega t$
$\ddot{Q}_{y}+2\gamma\dot{Q}_{y}+\Omega_{0}^{2}Q_{y}+\beta Q_{y}\left(Q_{x}^{2}+Q_{y}^{2}\right)$ $=Z_{y}E_{y}\sin\omega t.$ (A 7)
We first derive the two-component analogue of Eq. (9), which is given by
$\ddot{Q}_{x/y,\mathrm{even}}+2\gamma\dot{Q}_{x/y,\mathrm{even}}$
$\qquad+\left[\tilde{\Omega}_{x/y}^{2}\pm\frac{\alpha}{2}\left(3F_{x/y,1}^{2}-F_{y/x,1}^{2}\right)\cos\left(2\omega t\right)\right]Q_{x/y,\mathrm{even}}$
$\qquad+\alpha F_{x/y,1}F_{y/x,1}\sin\left(2\omega t\right)Q_{y/x,\mathrm{even}}=0$ (A 8)
with $\tilde{\Omega}_{x/y}^{2}=\Omega_{0}^{2}+\alpha\left(3F_{x/y}^{2}+F_{y/x}^{2}\right)/2$, where $F_{y}$ is defined analogously to $F_{x}$ in Eq. (7). As above, we expect parametric resonances near $2\omega=\tilde{\Omega}_{x/y}$, where $\tilde{\Omega}_{x}\neq\tilde{\Omega}_{y}$ for $F_{x}\neq F_{y}$. For now, let us choose the case $2\omega=\tilde{\Omega}_{x}$, such that the instability occurs for the $Q_{x}\left(t\right)$ component.
The oscillating terms in Eqs. (A 8) couple harmonics with frequencies $2\omega$, $4\omega$, … and DC terms. For the stability analysis, we therefore choose the ansatz
$Q_{i,\mathrm{even}}$ $=a_{i,1}\sin\left(2\omega t\right)+a_{i,2}\sin\left(4\omega t\right)+b_{i,0}$
$+b_{i,1}\cos\left(2\omega t\right)+b_{i,2}\cos\left(4\omega t\right).$ (A 9)
Furthermore, we neglect $\gamma$ for the duration of this analysis. While $\gamma$ determines the instability threshold amplitudes of the electromagnetic fields [see Eq. (12)], its effects become less important for driving amplitudes above the threshold, i.e., for any driving amplitude, $\gamma$ can be always chosen small enough that our analysis is accurate. We again search for the instability window for the detuning $\Delta=2\omega-\tilde{\Omega}_{x}$, such that the mode amplitudes $a_{i,n}$ and $b_{i,n}$ grow exponentially for $\Delta_{\mathrm{min}}<\Delta<\Delta_{\mathrm{max}}$. At $\Delta=\Delta_{\mathrm{max}/\mathrm{min}}$, the amplitudes will be constant. The boundaries of the instability interval $\Delta_{\mathrm{max}/\mathrm{min}}$ are then found by inserting the ansatz (A 9) into Eqs. (A 8) and assuming that $a_{i,n}$ and $b_{i,n}$ are indeed constant (Landau_Lifshitz_Mechanics, ). After a lengthy calculation, in which we compare the coefficients of different harmonics after inserting the ansatz (A 9) into Eqs. (A 8), we find that, to fourth order in $F_{x,1}$ and $F_{y,1}$
$\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ $=F_{y,1}^{4}\left(\frac{\beta^{2}}{16\Omega_{0}^{3}}+\frac{287\beta^{4}F_{x,1}^{4}}{576\Omega_{0}^{7}}+\frac{47\beta^{3}F_{x,1}^{2}}{192\Omega_{0}^{5}}\right)$
$+F_{y,1}^{2}\left(\frac{77\beta^{3}F_{x,1}^{4}}{192\Omega_{0}^{5}}-\frac{5\beta^{2}F_{x,1}^{2}}{8\Omega_{0}^{3}}\right)+\frac{9\beta^{2}F_{x,1}^{4}}{16\Omega_{0}^{3}}.$ (A 10)
The full result is too long to be quoted here but is easily found using computer algebra. Eq. (A 10) is valid for small $F_{x}$ and $F_{y}$, i.e. for small driving amplitudes. It is interesting to study the behavior of $\Delta$ close to $F_{y,1}=F_{x,1}$, i.e. for nearly perfect circular polarization. Writing $F_{y,1}=F_{x,1}+F_{\epsilon}$, we find
$\Delta_{\mathrm{max}}-\Delta_{\mathrm{min}}$ $=-F_{\epsilon}^{3}\left(\frac{9\Omega_{0}}{8F_{x,1}^{3}}+\frac{\beta^{2}F_{x,1}}{\Omega_{0}^{3}}+\frac{51\beta}{16F_{x,1}\Omega_{0}}\right).$ (A 11)
Notice that for $F_{y,1}\geq F_{x,1}$ (we choose both amplitudes positive w.l.o.g.), we have $\Delta_{\mathrm{max}}\leq\Delta_{\mathrm{min}}$, which indicates that the system is stable. For $F_{y,1}>F_{x,1}$, the $Q_{x}$ and $Q_{y}$ components switch places, and the instability occurs for $2\omega=\tilde{\Omega}_{y}$. The analysis for this case is completely analogous with $F_{x,1}$ and $F_{y,1}$ , as well as $\Omega_{x}$ and $\Omega_{y}$ interchanged. We therefore conclude that, in general, the instability occurs either for the $Q_{x}$ or the $Q_{y}$ component, depending on whether $F_{x,1}$ or $F_{y,1}$ is larger. For $E_{x}=E_{y}$, resulting in $F_{x}=F_{y}$, we find $\Delta_{\mathrm{min}}=\Delta_{\mathrm{max}}$ and the instability cannot be reached. (Note2, ).